OpenStudy (anonymous):

Which ordered pair is a solution of the equation y = -7x + 2 (1,2) (8,-54) (5, -35) (1,-7)

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

(8, -54) plug in -54= -7(8) +2 -54=-56+2 ? -54=-54